# A computable version of Hall’s Harem Theorem and Geometric von Neumann Conjecture

### v3

@@ 5.9. Remarks before part 2
Before the rest of the step $n+1$, we describe the goals which we want to achieve after the step:
@@
    the graph $\Gamma^{(n+1)}$ obtained at the end of the step is $U^{(n+1)}$-reflected and the corresponding graph $\Gamma^{(n+1)\star}$ satisfies Hall’s $d$-harem condition.
@@
    there is a sequence of vertices $u^{0}_{n},u^{1}_{n},u^{2}_{n},\ldots,u^{j}_{n}$, $1\leq j\leq n$ such that each edge $(u^{i}_{n},v_{u^{i-1}_{n}})$ belongs to $M_{n}$, and for some $10\leq\ell\leq j-$ 1 the sequence $(u^{\ell}_{n},u^{\ell+1}_{n},\ldots,u^{j}_{n})$ is a cycle;
@@
    there is a sequence of vertices $u^{0}_{n},u^{1}_{n},u^{2}_{n},\ldots,u^{j}_{n}$, $j\leq n$ such that each edge $(u^{i}_{n},v_{u^{i-1}_{n}})$ belongs to $M_{n}$, and $v_{u^{j}_{n}}$ is already adjacent to some edge from $\bigcup\limits_{i=0}^{n-1}M_{i}$;
@@ 5.10. Step n+1, part 2
    $v_{u^{l}_{n}}\in\Gamma^{(n)}$, but $v_{u^{l}_{n}}\notin\Gamma^{(n)}(-u^{0}_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$ (this case is impossible for $l=0$);
@@
    $v_{u^{l}_{n}}\in\Gamma^{(n)\star}$.
In case (1) $v_{u^{l}_{n}}$ was already added to $M$ at preceding steps and condition (ii) of Remarks at point 5.9 is satisfied. We finish part 2 of step $n+1$ without a new cycle.
In case (2) the last iteration closes the cycle $(u^{k}_{n},\ldots,u^{l}_{n})$. The length of this cycle is not greater than $l+1$.
@@
Case (3) is the most complicated case. We finish this part of the step by taking a new term $u^{l+1}_{n}:=u_{v^{l}_{n,1}}$. We force a cycle $(u^{l}_{n},u^{l+1}_{n})$ of length $2$ in part 3 of the step.
@@ 5.11. Example
Before we move to the third part of the step, we give an example of a cycle obtained by construction in case (2) of this part of the step.
The following pictures show how a cycle of length $3$ can be obtained by this procedure in the graph satisfying Hall’s $3$-harem condition.
@@
Figure 9. The final $M_{n}$ with the obtained cycle of length $3$ marked in blue. The red edge is the edge from $\mathfrak{M}^{n}_{1}$ that was not added to $M_{n}^{0}$.
@@ 5.12. Step n+1, part 3
In the last case the edge $(u^{l}_{n},v^{l}_{n,1})$ is in $\Gamma^{(n)}$ and $v_{u^{l}_{n}}$ is not in $\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$, thus applying $U^{(n)}$-reflectedness of $\Gamma^{(n)}$ we see $(u^{l+1}_{n},v_{u^{l}_{n}})\in\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$. Observe that since $u^{\perp}_{j_{1}},\ldots,u^{\perp}_{j_{l}}$ are in $U^{(n)\perp}$,
@@
By part 1 of this step the graph $\Gamma^{(n)\star}(-u_{n})$ satisfies $c.e.H.h.c.(d)$, thus we can apply Proposition 5.1 for $\Gamma^{(n)\star}(-u_{n})$ (viewed as $\Gamma^{\prime}$ in formulation, with $\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$ viewed as $\Gamma$) and compute $\mathfrak{M}^{2}_{n}$, a perfect $(1,d)$-matching in the ball $\mathcal{B}^{(n)\star}(u^{l+1}_{n})$.
We now check whether there is $j$ with $u^{l+1}_{n}=u^{\perp}_{j}$. If it exists, we set $\dot{v}^{l+1}_{n,i}:=v^{\perp}_{j,i}$ for $11\leq i\leq d-$ 1. When it does not exist, $\dot{v}^{l+1}_{n,1},\ldots,\dot{v}^{l+1}_{n,d}$ will denote the elements adjacent to $u^{l+1}_{n}$ under $\mathfrak{M}^{2}_{n}$.
@@
In either case we produce a cycle of length 2 by including the pair $(u^{l+1}_{n},v_{u^{l}_{n}})$ into $M^{l+1}_{n}$. In fact we include it into $M^{l+1}_{n}$ together with a fan with the root $u^{l+1}_{n}$ and $(d-2)$ leaves taken among $\dot{v}^{l+1}_{n,i}$.
To be precise we organize it as follows.
In case A) If $k=d$, we set $v^{l+1}_{n,i-1}:=\dot{v}^{l+1}_{n,i}$ for $2\leq i\leq d$. If $k\neq d$ we set $v^{l+1}_{n,i}:=\dot{v}^{l+1}_{n,i}$ for $11\leq i\leq d-$ 1. The set $M^{l+1}_{n}$ consists of edges $(u^{l+1}_{n},v^{l+1}_{n,i})$ for $11\leq i\leq d-$ 1. The procedure is finished.
In case B) $v^{l+1}_{n,1}=v_{u^{l}_{n}}$ and $v^{l+1}_{n,i+1}=\dot{v}^{l+1}_{n,i}$, $21\leq i\leq d-$ 2.333It is worth noting here that if $u^{l+1}_{n}$ coincides with some $u^{\perp}_{j}$, then the vertex $\dot{v}^{l+1}_{n,d}$ does not exist, i.e the only vertex from the fan of $u^{l+1}_{n}$ that is outside of $M^{l+1}_{n}$ is $\dot{v}^{l+1}_{n,d-1}$. We define $\dot{u}_{n}^{\perp}:=u$ and rename the remaining $d-1$ vertices $v_{j}$ to $\dot{v}^{\perp}_{n,i}$. The procedure is finished.
@@ 5.13. Output of the step
This is $M_{n}$ and $\Gamma^{(n+1)}$. It remains to produce graphs $\Gamma^{(n+1)\star},\Gamma^{(n+1)\perp}$. To do that we define auxilliary graphs $\mathfrak{T}$ and $\dot{\Gamma}^{(n)\perp}$.
In cases (1), (2) and (3)A) we set $\mathfrak{U}:=U^{(n+1)}\setminus U^{(n)\perp},\mathfrak{V}:=V^{(n+1)}\setminus V^{(n)\perp}$. Let $\mathfrak{T}=(\mathfrak{U},\mathfrak{V})$ and $\dot{\Gamma}^{(n)\perp}:=\Gamma^{(n)\perp}\cap\Gamma^{(n+1)}$. Note here that in cases (1) and (2) only elements of $\Gamma^{(n)\perp}$ were added to $M_{n}$ at the part 2 of the step. Therefore in these cases $\mathfrak{T}$ coincides with $\Gamma^{(n)\star}(-u_{n})$ and satisfies Hall’s $d$-harem condition.
In case (3)B) we set $\mathfrak{U}:=U^{(n+1)}\setminus(U^{(n)\perp}\cup\{\dot{u}_{n}^{\perp}\}),\mathfrak{V}:=V^{(n+1)}\setminus(V^{(n)\perp}\cup\{\dot{v}_{n,i}^{\perp}:1\leq i\leq d-1\})$. Let $\mathfrak{T}=(\mathfrak{U},\mathfrak{V})$ and $\dot{U}^{(n)\perp}:=(U^{(n)\perp}\cup\{\dot{u}_{n}^{\perp}\})\cap U^{(n+1)}$. We define $\dot{V}^{(n)\perp}$ accordingly: $\dot{V}^{(n)\perp}:=(V^{(n)\perp}\cup\{\dot{v}_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}$, and set
$\dot{\Gamma}^{(n)\perp}:=(\dot{U}^{(n)\perp},\dot{V}^{(n)\perp}).$
@@
In the case (3)B) (i.e. $\dot{u}_{n}^{\perp}$ exists) if $u_{i}^{\perp}\neq\dot{u}_{n}^{\perp}\neq u_{j}^{\perp}$ (for any output of the claim), we set $u_{n}^{\perp}:=\dot{u}_{n}^{\perp}$ and $v_{n,k}^{\perp}:=\dot{v}_{n,k}^{\perp}$ for $11\leq k\leq d-$ 1. Otherwise, or in the remaining cases $u_{n}^{\perp},v_{n,k}^{\perp}$ are not defined.
@@
$V^{(n+1)\perp}:=(V^{(n)\perp}\cup\{v_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}.$
@@
$V^{(n+1)\perp}:=((V^{(n)\perp}\setminus\{v_{j,k}^{\perp}:1\leq k\leq d-1\})\cup\{v_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}.$
@@
$V^{(n+1)\perp}:=((V^{(n)\perp}\setminus\{v_{i,k}^{\perp},v_{j,k}^{\perp}:1\leq k\leq d-1\})\cup\{v_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}.$
6. Technical Lemmas
The notation used in this section does not correspond to the construction above.
@@ 6. Technical Lemmas
Let $\Gamma=(U,V,E)$ be $U$-reflected and let $\Gamma^{\star}$ be a subgraph of $\Gamma$ induced by the sets of vertices $U^{\star}=U\setminus U^{\perp}$, $V^{\star}=V\setminus V^{\perp}$. Assume that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition and assume that for each $Y\subset U^{\perp}$ we have $|N_{\Gamma^{\perp}}(Y)|\geq(d-1)|Y|$.
@@
Let $U_{X}:=\{u\in U:v_{u}\in X\}$. By $U$-reflectedness of $\Gamma$ we have $|U_{X}|=|X|$. Consider the sets
@@
Since $d\geq 2$ it follows that
@@
This lemma cannot be applied in the situation of Part 1 of the step of the main construction, since some vertices from a $U$-reflected graph are removed. In that case, we have a graph $\Gamma$ which is not $U$-reflected, but for some $v\in V$ its subgraph $\Gamma(-v)$ is $U$-reflected. Because of this, in the following lemmas both $\Gamma^{\perp}$ and $\Gamma^{\star}$ are the subgraphs of $\Gamma(-v)$.
@@
Assume that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition and assume that for each $Y\subset U^{\perp}$ we have $|N_{\Gamma}(Y)\cap V^{\perp}|\geq(d-1)|Y|$.
@@
Moreover, the equality $|N_{\Gamma}(X)|=(d-1)|X|-1$ can happen only if $v\in N_{\Gamma}(U_{X})$ and $N_{\Gamma^{\star}}(U_{X})=\emptyset$, where $U_{X}:=\{u\in U:v_{u}\in X\}$.
@@
$U_{X}^{\star}:=U_{X}\setminus U^{\perp}\,\mbox{ and }\,U_{X}^{\perp}:=U_{X}\cap U^{\perp}.$
@@
Observe that $|N_{\Gamma(-v)}(U_{X})|=(d-1)|U_{X}|$ only if $N_{\Gamma(-v)}(U_{X})\subseteq V^{\perp}$, i.e. $N_{\Gamma^{\star}}(U_{X})=\emptyset$. If $v\notin N_{\Gamma}(U_{X})$ then
@@
Another fact that will be regularly used in the proofs of Lemmas 7.1 and 7.5 is the following.
@@
  • $M_{u}$ be a perfect $(1,d)$-matching in the ball $\mathcal{B}(u)$;
  • $U_{1}$ be a subset of $U$ from the ball of $u$ of radius $\leq 2$, $|U_{1}|\leq k$, and $V_{1}$ be the set of vertices adjacent in $M_{u}$ to vertices from $U_{1}$.
@@
Let us show that the inequality $|N_{\Gamma^{\prime}}(X)|-d|X|\geq 0$ holds for any $X\subset U\setminus U_{1}$. If $X$ is not connected, then its neighbourhood is a disjoint union of the neighbourhoods of the connected components of $X$. Thus the inequality for each component implies it for $X$. Therefore without loss of generality we may also assume that $X$ is connected. Further, we may assume that $X\nsubseteq U\cap\mathcal{B}(u)$ and $V_{1}\cap N_{\Gamma}(X)\neq\emptyset$. Indeed, if $V_{1}\cap N_{\Gamma}(X)=\emptyset$ then $N_{\Gamma^{\prime}}(X)=N_{\Gamma}(X)$, i.e. $|N_{\Gamma^{\prime}}(X)|\geq d|X|$. On the other hand if $X\subset U\cap\mathcal{B}(u)$ then using existence of an $(1,d)$-matching $M_{u}$ from $U\cap\mathcal{B}(u)$ to $V\cap\mathcal{B}(u)$ which is perfect for subsets of $U\cap\mathcal{B}(u)$, we see that $|N_{\Gamma^{\prime}}(X)|-d|X|\geq 0$.
Let us choose $u_{1},u_{2}\in X$ such that $u_{1}\in N_{\Gamma}(V_{1})$ and $u_{2}\in X\setminus\mathcal{B}(u)$. By the choice of the radius of the ball $\mathcal{B}(u)$ the distance between $u_{1}$ and $u_{2}$ is at least $12h(2d)+$ 1. Thus $|X|\geq h(kd)$. Since $\Gamma$ satisfies $c.e.H.h.c.(d)$ we have
@@
Let $\Gamma:=(U,V,E)$ be a bipartite graph, and $\Gamma^{\star}=(U^{\star},V^{\star},E^{\star})$ be a subgraph in $\Gamma$ satisfying Hall’s $d$-harem condition. Assume that $M$ is a corresponding perfect $(1,d)$-matching in $\Gamma^{\star}$.
@@
In the technical lemma below we use the following notion of Kierstead, [Kie83]. The subset $X$ of $U$ (resp. of $V$) is called connected if for all $x,x^{\prime}\in X$ there exist a path $x=p_{0},p_{1},\ldots,p_{k}=x^{\prime}$ in $\Gamma$ such that $p_{i}\in X\cup N_{\Gamma}(X)$ for all $i\leq k$. Note here that this definition concerns only bipartite graphs.
@@
Let $\Gamma=(U,V,E)$ be a bipartite graph. Let $\Gamma^{\star}=(U^{\star},V^{\star},E^{\star})$ be a subgraph of $\Gamma$ satisfying Hall’s $d$-harem condition and $M$ be a corresponding perfect $(1,d)$-matching.
@@
  • $\widehat{u}\in N_{\Gamma}(X)\setminus U^{\star}$;
@@
It is clear that $\widehat{v}\in X$. Let $X^{\prime}$ denote the subset of all elements of $X$ that are accessible from $\widehat{u}$ through $X$ by the matching $M$. Assume that $\widehat{v}\not\in X^{\prime}$. Since $\widehat{u}\in N_{\Gamma}(X)$, we then see that there exist $u^{\prime}_{0},v^{\prime}_{0}$ such that $\widehat{u},u^{\prime}_{0}\in N_{\Gamma}(v^{\prime}_{0})$ and $(u^{\prime}_{0},v^{\prime}_{0})\in M$. Therefore $X^{\prime}\neq\emptyset$. We will show that $|N_{\Gamma}(X\setminus X^{\prime})|<\frac{1}{d}|X\setminus X^{\prime}|$.
@@
$|N_{\Gamma}(X\setminus X^{\prime})|\leq N_{\Gamma}(X)|-|U_{M}(X^{\prime})|\leq|N_{\Gamma}(X)|-\lceil\frac{l}{d}\rceil<\frac{1}{d}(|X|-l)=\frac{1}{d}|X\setminus X^{\prime}|.$
@@
This lemma will be essential in the proofs below in order to replace some $(1,d-1)$-fans in the matching. Since in our case this matching is partial and finite we need an additional technical lemma. It will guarantee that in a graph satisfying $c.e.H.h.c.(d)$ Hall’s $d$-harem condition can be preserved after some replacement of $(1,d-1)$-fans.

### v4

@@ 5.9. Remarks before part 2
Before the rest of step $n+1$, we describe the goals which we want to achieve after the step:
@@
    the graph $\Gamma^{(n+1)}$ obtained at the end of the step, is $U^{(n+1)}$-reflected, and the corresponding graph $\Gamma^{(n+1)\star}$ satisfies Hall’s $d$-harem condition.
@@
    there is a sequence of vertices $u^{0}_{n},u^{1}_{n},u^{2}_{n},\ldots,u^{j}_{n}$, $1\leq j\leq n$, such that every edge $(u^{i}_{n},v_{u^{i-1}_{n}})$ belongs to $M_{n}$, and for some $0\leq\ell\leq j-1$ the sequence $(u^{\ell}_{n},u^{\ell+1}_{n},\ldots,u^{j}_{n})$ is a cycle;
@@
    there is a sequence of vertices $u^{0}_{n},u^{1}_{n},u^{2}_{n},\ldots,u^{j}_{n}$, $j\leq n$, such that each edge $(u^{i}_{n},v_{u^{i-1}_{n}})$ belongs to $M_{n}$, and $v_{u^{j}_{n}}$ is already adjacent to some edge from $\bigcup\limits_{i=0}^{n-1}M_{i}$;
@@ 5.10. Step n+1, part 2
    $v_{u^{l}_{n}}\in\Gamma^{(n)\perp}$, but $v_{u^{l}_{n}}\notin\Gamma^{(n)}(-u^{0}_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$ (this case is impossible for $l=0$);
@@
    $v_{u^{l}_{n}}\in\Gamma^{(n)\star}$ (and obviously $v_{u^{l}_{n}}\in\Gamma^{(n)}(-u^{0}_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$).
In case (1), $v_{u^{l}_{n}}$ was already added to $M$ at preceding steps and condition (ii) of Remarks at point 5.9 is satisfied. We finish part 2 of step $n+1$ without a new cycle.
In case (2), the last iteration closes the cycle $(u^{k}_{n},\ldots,u^{l}_{n})$. The length of this cycle is not greater than $l+1$.
@@
Case (3) is the most complicated one; part 3 of the step will be entirely dedicated to it. In this case, we finish part 2 of the step by taking a new term $u^{l+1}_{n}:=u_{v^{l}_{n,1}}$. In part 3 we will force the cycle $(u^{l}_{n},u^{l+1}_{n})$ of length $2$.
@@ 5.11. Example
Before we move to the third part of the step, we give an example of a cycle obtained by the algorithm in case (2) of this part of the step.
The following pictures show how a cycle of length $3$ arises in this procedure when the graph satisfies Hall’s $3$-harem condition.
@@
Figure 10. The final $M_{n}$ with the obtained cycle of length 33 marked in blue. The red edge is the edge from $\mathfrak{M}^{n}_{1}$ that was not added to $M_{n}^{0}$.
@@ 5.12. Step n+1, part 3
Case (3) of part 2 guarantees that the edge $(u^{l}_{n},v^{l}_{n,1})$ is in $\Gamma^{(n)}$ and $v_{u^{l}_{n}}\in\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$. Thus applying $U^{(n)}$-reflectedness of $\Gamma^{(n)}$ we see $(u^{l+1}_{n},v_{u^{l}_{n}})\in\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$. Observe that since in part 2 the elements $u^{\perp}_{j_{1}},\ldots,u^{\perp}_{j_{l}}$ were chosen in $U^{(n)\perp}$,
@@
By part 1 of this step the graph $\Gamma^{(n)\star}(-u_{n})$ satisfies $c.e.H.h.c.(d)$, thus we can apply Proposition 5.1 for $\Gamma^{(n)\star}(-u_{n})$ (viewed as $\Gamma^{\prime}$ in the formulation, with $\Gamma^{(n)}(-u_{n},-u^{1}_{n},\ldots,-u^{l}_{n})$ viewed as $\Gamma$) and compute $\mathfrak{M}^{2}_{n}$, a perfect $(1,d)$-matching in the ball $\mathcal{B}^{(n)\star}(u^{l+1}_{n})$.
We now check whether there is $j$ with $u^{l+1}_{n}=u^{\perp}_{j}$. If it exists, we set $\dot{v}^{l+1}_{n,i}:=v^{\perp}_{j,i}$ for $1\leq i\leq d-1$. When it does not exist, $\dot{v}^{l+1}_{n,1},\ldots,\dot{v}^{l+1}_{n,d}$ will denote the elements adjacent to $u^{l+1}_{n}$ under $\mathfrak{M}^{2}_{n}$.
@@
In either case we produce a cycle of length 2 by including the pair $(u^{l+1}_{n},v_{u^{l}_{n}})$ into $M^{l+1}_{n}$. In fact we include it into $M^{l+1}_{n}$ together with a fan with the root $u^{l+1}_{n}$ and (d2)(d-2) leaves taken among $\dot{v}^{l+1}_{n,i}$.
To be precise, we organize it as follows.
In case A) if $k=d$, we set $v^{l+1}_{n,i-1}:=\dot{v}^{l+1}_{n,i}$ for $2\leq i\leq d$. If $k\neq d$ we set $v^{l+1}_{n,i}:=\dot{v}^{l+1}_{n,i}$ for $1\leq i\leq d-1$. The set $M^{l+1}_{n}$ consists of edges $(u^{l+1}_{n},v^{l+1}_{n,i})$ for $1\leq i\leq d-1$. The procedure is finished.
In case B) $v^{l+1}_{n,1}=v_{u^{l}_{n}}$ and $v^{l+1}_{n,i+1}=\dot{v}^{l+1}_{n,i}$, $1\leq i\leq d-2$. The set $M^{l+1}_{n}$ consists of edges $(u^{l+1}_{n},v^{l+1}_{n,i})$ for $1\leq i\leq d-1$. We define $\dot{u}_{n}^{\perp}:=u$ and rename the remaining d1d-1 vertices $v_{j}$ to $\dot{v}^{\perp}_{n,i}$. The procedure is finished.
@@ 5.13. Output of the step
We have already defined $M_{n}$ and $\Gamma^{(n+1)}$. It remains to produce graphs $\Gamma^{(n+1)\star},\Gamma^{(n+1)\perp}$. For this purpose, we define auxiliary graphs
$\mathfrak{T}=(\mathfrak{U},\mathfrak{V})\mbox{ and }\dot{\Gamma}^{(n)\perp}=(\dot{U}^{(n)\perp},\dot{V}^{(n)\perp}).$
In cases (1), (2) and (3)A) of parts 2 and 3 we set $\mathfrak{U}:=U^{(n+1)}\setminus U^{(n)\perp}$, $\mathfrak{V}:=V^{(n+1)}\setminus V^{(n)\perp}$, and $\dot{\Gamma}^{(n)\perp}:=\Gamma^{(n)\perp}\cap\Gamma^{(n+1)}$. Note here that in cases (1) and (2), in part 2 of the step, only elements of $\Gamma^{(n)\perp}$ were added to $M_{n}$. Therefore, in these cases, $\mathfrak{T}$ coincides with $\Gamma^{(n)\star}(-u_{n})$ and satisfies Hall’s $d$-harem condition.
In case (3)B) we set
  • $\mathfrak{U}:=U^{(n+1)}\setminus(U^{(n)\perp}\cup\{\dot{u}_{n}^{\perp}\})$,
  • $\mathfrak{V}:=V^{(n+1)}\setminus(V^{(n)\perp}\cup\{\dot{v}_{n,i}^{\perp}:1\leq i\leq d-1\})$,
unless $u^{l+1}_{n}$ coincides with some $u^{\perp}_{j}$. In the latter case, $\dot{v}^{l+1}_{n,d-1}$ is the only vertex left from the fan of $u^{\perp}_{j}$ in $\Gamma^{(n)\perp}$.  [^3] Furthermore, the vertex $\dot{v}^{l+1}_{n,d}$ does not exist, i.e the only vertex from the fan of $u^{l+1}_{n}$ that is outside of $M^{l+1}_{n}$ is $\dot{v}^{l+1}_{n,d-1}$. Consequently, we define
$\bullet\,\mathfrak{V}:=(V^{(n+1)}\cup\{\dot{v}^{l+1}_{n,d-1}\})\setminus(V^{(n)\perp}\cup\{\dot{v}_{n,i}^{\perp}:1\leq i\leq d-1\})$.
We define $\dot{U}^{(n)\perp}$ and $\dot{V}^{(n)\perp}$ accordingly:
  • $\dot{U}^{(n)\perp}:=(U^{(n)\perp}\cup\{\dot{u}_{n}^{\perp}\})\cap U^{(n+1)}$,
  • $\dot{V}^{(n)\perp}:=(V^{(n)\perp}\cup\{\dot{v}_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}$,
@@
In case (3)B) (i.e. $\dot{u}_{n}^{\perp}$ exists) if $u_{i}^{\perp}\neq\dot{u}_{n}^{\perp}\neq u_{j}^{\perp}$ (for any output of the claim), we set $u_{n}^{\perp}:=\dot{u}_{n}^{\perp}$ and $v_{n,k}^{\perp}:=\dot{v}_{n,k}^{\perp}$ for $1\leq k\leq d-1$. Otherwise, or in the remaining cases $u_{n}^{\perp},v_{n,k}^{\perp}$ are not defined.
@@
  • $V^{(n+1)\perp}:=((V^{(n)\perp}\setminus\{\dot{v}^{l+1}_{n,d-1}\})\cup\{v_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}.$
@@
  • $V^{(n+1)\perp}:=(((V^{(n)\perp}\setminus\{\dot{v}^{l+1}_{n,d-1}\})\setminus\{v_{j,k}^{\perp}:1\leq k\leq d-1\})\cup\{v_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}.$
@@
  • $V^{(n+1)\perp}:=(((V^{(n)\perp}\setminus\{\dot{v}^{l+1}_{n,d-1}\})\setminus\{v_{i,k}^{\perp},v_{j,k}^{\perp}:1\leq k\leq d-1\})\cup\{v_{n,k}^{\perp}:1\leq k\leq d-1\})\cap V^{(n+1)}.$
6. General Lemmas
The construction presented in Section 5 is supported by the lemmas of Section 7. In order to prove them, we need some additional observations of slightly more general character. This is the purpose of this section. We warn the reader that the notation used here does not coincide with that of Section 5.
@@ 6. General Lemmas
Let $\Gamma=(U,V,E)$ be $U$-reflected and let $\Gamma^{\star}$ be a subgraph of $\Gamma$ induced by the sets of vertices $U^{\star}=U\setminus U^{\perp}$, $V^{\star}=V\setminus V^{\perp}$. Assume that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition, and assume that for each $Y\subset U^{\perp}$ we have $|N_{\Gamma}(Y)\cap V^{\perp}|\geq(d-1)|Y|$.
@@
Let $U_{X}:=\{u\in U:v_{u}\in X\}$. Since $\Gamma$ is $U$-reflected, $|U_{X}|=|X|$. Consider the sets
@@
Since $d\geq 2$, it follows that
@@
The conditions of this lemma do not fit to the situation of Part 1 of the step of the main construction. Indeed, when some vertices from the graph are removed, we cannot guarantee that it is still $U$-reflected. On the other hand, it can happen that for some $v\in V$, the subgraph $\Gamma(-v)$ is $U$-reflected. This is the reason why in the following lemmas both $\Gamma^{\perp}$ and $\Gamma^{\star}$ are subgraphs of some $\Gamma(-v)$.
@@
Assume that $\Gamma^{\star}$ satisfies Hall’s $d$-harem condition, and assume that for each $Y\subset U^{\perp}$ we have $|N_{\Gamma}(Y)\cap V^{\perp}|\geq(d-1)|Y|$.
@@
Moreover, the equality $|N_{\Gamma}(X)|=(d-1)|X|-1$ can happen only if $v\in N_{\Gamma}(U_{X})$ and $N_{\Gamma^{\star}}(U_{X})=\emptyset$, where $U_{X}=\{u\in U:v_{u}\in X\}$.
@@
[…]
@@
Observe that $|N_{\Gamma(-v)}(U_{X})|=(d-1)|U_{X}|$ only if $N_{\Gamma(-v)}(U_{X})\subseteq V^{\perp}$, i.e. $N_{\Gamma^{\star}}(U_{X})=\emptyset$ (since Hall’s $d$-harem condition is satisfied for any subset of $U^{\star}$). If $v\notin N_{\Gamma}(U_{X})$, then
@@
The following proposition will be used regularly in the proofs of Lemmas 7.1 and 7.5.
@@
Assume that $\Gamma=(U,V,E)$ is a highly computable bipartite graph satisfying $c.e.H.h.c.(d)$ with witness $h$. Let:
@@
  • $M_{u}$ be a $(1,d)$-matching in $\mathcal{B}(u)$ which satisfies the conditions of the perfect $(1,d)$-matchings for all vertices that are at the distance less than $\max\{2h(kd)+3,5\}$ from $u$;
  • $U_{1}\subseteq U$ be a subset of the ball of $u$ of radius $\leq 2$, $|U_{1}|\leq k$, and $V_{1}$ be the set of vertices adjacent in $M_{u}$ to vertices from $U_{1}$.
@@
Let us show that the inequality $|N_{\Gamma^{\prime}}(X)|-d|X|\geq 0$ holds for any $X\subset U\setminus U_{1}$. If $X$ is not connected (in the sense of the Definition 2.1), then its neighborhood is a disjoint union of the neighborhoods of the connected components of $X$. Thus the inequality for each component implies it for $X$. Therefore, without loss of generality we may also assume that $X$ is connected. Further, we may assume that $X\nsubseteq U\cap\mathcal{B}(u)$ and $V_{1}\cap N_{\Gamma}(X)\neq\emptyset$. Indeed, if $V_{1}\cap N_{\Gamma}(X)=\emptyset$ then $N_{\Gamma^{\prime}}(X)=N_{\Gamma}(X)$, i.e. $|N_{\Gamma^{\prime}}(X)|\geq d|X|$. On the other hand, if $X\subset U\cap\mathcal{B}(u)$ then using existence of an $(1,d)$-matching $M_{u}$ from $U\cap\mathcal{B}(u)$ to $V\cap\mathcal{B}(u)$ which is perfect for subsets of $U\cap\mathcal{B}(u)$, we see that $|N_{\Gamma^{\prime}}(X)|-d|X|\geq 0$.
Let us choose $u_{1},u_{2}\in X$ such that $u_{1}\in N_{\Gamma}(V_{1})$ and $u_{2}\in X\setminus\mathcal{B}(u)$. By the choice of the radius of the ball $\mathcal{B}(u)$ the distance between $u_{1}$ and $u_{2}$ is at least $2h(kd)+1$. Thus $|X|\geq h(kd)$. Since $\Gamma$ satisfies $c.e.H.h.c.(d)$, we have
@@
Let $\Gamma:=(U,V,E)$ be a bipartite graph. Assume that $\Gamma^{\star}=(U^{\star},V^{\star},E^{\star})$ is a subgraph of $\Gamma$ satisfying Hall’s $d$-harem condition, and $M$ is a perfect $(1,d)$-matching in $\Gamma^{\star}$.
@@
Figure 12. $x$ is accessible from $y$ by matching $M$ (red), the elements of $X$ are blue and the elements of $N_{\Gamma}(X)$ are green
@@
Let $\Gamma=(U,V,E)$ be a bipartite graph. Assume that $\Gamma^{\star}=(U^{\star},V^{\star},E^{\star})$ is a subgraph of $\Gamma$ satisfying Hall’s $d$-harem condition, and $M$ is a perfect $(1,d)$-matching in $\Gamma^{\star}$.
@@
  • $\widehat{u}\in N_{\Gamma}(X)\setminus U^{\star}$.
  • $(\widehat{u},\widehat{v})\notin E$.
@@
It is clear that $\widehat{v}\in X$. Let $X^{\prime}$ denote the subset of all elements of $X$ that are either adjacent to $\widehat{u}$ or accessible from $\widehat{u}$ through $X$ by the matching $M$. Since $\widehat{u}\in N_{\Gamma}(X)$, $X^{\prime}\neq\emptyset$. In order to get a contradiction, assume that $\widehat{v}\not\in X^{\prime}$. We will show that $|N_{\Gamma}(X\setminus X^{\prime})|<\frac{1}{d}|X\setminus X^{\prime}|$.
@@
$|N_{\Gamma}(X\setminus X^{\prime})|\leq|N_{\Gamma}(X)|-|U_{M}(X^{\prime})|\leq|N_{\Gamma}(X)|-\lceil\frac{l}{d}\rceil<\frac{1}{d}(|X|-l)=\frac{1}{d}|X\setminus X^{\prime}|.$
@@
This lemma will be used in situations when some (1,d1)(1,d-1)-fans in the matching are replaced by some new fans. The following technical lemma will additionally guarantee that Hall’s $d$-harem condition can be preserved after some replacement of (1,d1)(1,d-1)-fans.